# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ I Introduction
We organize the paper as follows. We begin by describing the model system in Section II. Next, Section III calculates the dipolar fields from the magnon-polaritons. A quantum-mechanical analysis of the magnon-photon coupling and the magnon-polariton dispersion and spin follows in Section IV. For comparison, Section V revisits this physics using a classical framework, detailing the polarization of the magnonic part. We conclude with a summary of our results in Section VI. Additional technical details are deferred to the appendices: the demagnetization tensor calculation (Appendix A), general dispersion relations (Appendix B), and the derivation of mode spin (Appendix C).
@@ II Model
III Dipolar fields
@@
where $\rho_{s}=\rho(1-(T/T_{c})^{4})$ and $\rho_{n}=\rho(T/T_{c})^{4}$ are superfluid and normal fluid densities, respectively. Equation (3) can then be rewritten as
@@
where $\sigma_{n}=\rho e^{2}\tau/m_{e}$ is the conductivity of normal metals, $\lambda_{L}=\sqrt{m_{e}/(\mu_{0}\rho_{s}e^{2})}$ is London’s penetration depth, $\rho$ is the electron density, $m_{e}$ is the electron mass, $e$ is the electron charge, $T_{c}$ is the critical temperature of the superconductor, $T$ is the temperature of the system, and $\tau$ is the relaxation time of electrons.
@@
On the other hand, after substituting explicit expression for $\lambda_{L}$ and $\sigma_{n}$ Eq. (4) is transformed to
@@
$\mathrm{in~S1:}~~\tilde{\bm{E}}(x)=\bm{S}_{1}e^{iB(x-d_{AF})},$
$\mathrm{in~S2:}~~\tilde{\bm{E}}(x)=\bm{S}_{2}e^{-iB(x+d_{AF})},$
@@
where $A=\sqrt{k_{AF}^{2}-k^{2}}$, $B=\sqrt{k_{S}^{2}-k^{2}}$, and $\bm{S}_{1(2)}=(S_{1(2)x},S_{1(2)y},S_{1(2)z})^{T}$ and $\bm{A}^{\pm}=(A^{\pm}_{x},A^{\pm}_{y},A^{\pm}_{z})^{T}$ are vectors of unknown coefficients. Here we have used the assumption that the AF is thin enough, such that the magnetizations $\bm{M}_{1,2}$ can be considered as constant along the film normal $x$-axis throughout $2d_{AF}$. Also, for the considered parameters $B\approx i/\lambda_{\rm{eff}}$, which will be substituted in all the expressions containing $B$ in the following text.
@@
This dipolar field is responsible for the formation of magnon-polaritons with ultrastrong coupling between magnon and photon, which we address below by the quantum (Sec. IV) and classical (Sec. V) views, respectively.
IV Magnon-polaritons: quantum view
@@
Then we introduce the operators of the spin density $\hat{\bm{S}}_{i}({\bm{r}})$ at the sublattice $i$, so that $\hat{\bm{M}}_{i}({\bm{r}})=-\gamma\hbar\hat{\bm{S}}_{i}({\bm{r}})$ and $M_{s}=\gamma\hbar S$, where $\gamma$ is the gyromagnetic ratio of electrons and $S$ is the total spin of the system. As the next step, we represent the spin-density operators $\hat{\bm{S}}_{i}({\bm{r}})$ in terms of the bosonic operators $\hat{a}_{i}({\bm{r}})$ via the Holstein-Primakoff transformation [40]:
@@
$+\frac{h_{-}}{\sqrt{h_{-}^{2}\left(\mu_{0}\gamma H_{0}+\sqrt{(\mu_{0}\gamma H_{0})^{2}+h_{-}^{2}}\right)^{2}}},$
@@
In the general case of arbitrary magnitudes of the applied magnetic fields, the eigenfrequencies of the magnon-polaritons are expressed by Eqs. (118) and are plotted in Fig. 2 for three typical values of the external magnetic field $H_{0}$: zero external field [Fig. 2(a)], moderate external field $\gamma\mu_{0}H_{0}\sim h_{-}$ [Fig. 2(b)], and strong external field $\gamma\mu_{0}H_{0}\gg h_{-}$ [Fig. 2(c)]. It is seen that at nonzero magnetic field both modes are coupled to the Swihart mode according to the fact that both coupling constants $g_{1}$ and $g_{2}$ are nonzero in this case. At large magnetic fields $\gamma\mu_{0}H_{0}\gg h_{-}$, both magnonic modes are equally coupled to the Swihart mode. In addition to being at the boundary of the ultra-strong coupling, the complex magnon-polariton excitations have other important properties. First of all, they are bosonic quasiparticles with non-integer average spin $\langle S_{z}\rangle$ smaller than $\hbar$. The average spin of all magnon-polariton excitation branches is shown by the color of the lines in Fig. 2. The details of the calculation of the average spin are presented in the Appendix C.
Figure 2: Dispersion of the eigenmodes in the S/AF/S structure in the cases of no external field [(a)], moderate external field $\mu_{0}H_{0}=0.05$ T [(b)], and strong external field $\mu_{0}H_{0}=0.5$ T [(c)]. The colors of the curves depict the value of the average spin $\langle S_{z}\rangle$ on the eigenmodes.
@@
V Magnon-polaritons: Classical view
Now we shall use an alternative classical approach to obtain the magnon-polariton eigenfrequencies and mode functions. The Landau-Lifshitz-Gilbert (LLG) equation for the magnetizations of the two sublattices $\bm{M}_{1,2}$ takes the form:
$\dfrac{\partial\bm{M}_{i}}{\partial t}=-\gamma\mu_{0}\bm{M}_{i}\times\bm{H}_{i},~~i\in\{1,2\},$ (60)
@@ V Magnon-polaritons: Classical view
is the total effective magnetic field acting on the magnetization of the sublattice $i$, $\overline{i}=2(1)$ if $i=1(2)$. Substituting Eq. (33) and a classical analogue of Eq. (36) into Eq. (61) and then linearizing Eq. (60) with respect to the magnon amplitudes, we obtain
$\left(\begin{array}[]{cccc}i\omega&0&-\gamma\mu_{0}(F_{+}+N_{y}M_{s})&-\gamma\mu_{0}(J_{AF}+N_{y}M_{s})\\ 0&i\omega&\gamma\mu_{0}(J_{AF}+N_{y}M_{s})&\gamma\mu_{0}(F_{-}+N_{y}M_{s})\\ \gamma\mu_{0}(F_{+}+N_{x}M_{s})&\gamma\mu_{0}(J_{AF}+N_{x}M_{s})&i\omega&0\\ -\gamma\mu_{0}(J_{AF}+N_{x}M_{s})&-\gamma\mu_{0}(F_{-}+N_{x}M_{s})&0&i\omega\end{array}\right)\left(\begin{array}[]{c}\tilde{M}_{1x}\\ \tilde{M}_{2x}\\ \tilde{M}_{1y}\\ \tilde{M}_{2y}\end{array}\right)=0,$ (70)
@@
are the eigenfrequencies in the case with no external magnetic field, $H_{0}=0$. As the demagnetization factor $N$ is expressed by Eq. (29) and depends on $\omega$ and $k$, Eq. (72) and Eq. (73b) are actually implicit equations for $\omega_{1,2}(k)$ and $\omega_{02}(k)$, respectively. Figure 3 shows the dispersion relations, calculated for the structure where the antiferromagnet is $\mathrm{MnF}_{2}$ with the same parameters addressed in Sec. IV. Here we again consider the bulk volume configuration, i.e., $\theta=0$. We note that according to the quantum formalism, the coupling constants between magnons and Swihart photon vanish in the Damon-Eshbach configuration $\theta=\pi/2$. The dispersion relations calculated via the quantum approach from Eq. (118) and via the classical approach from Eq. (72) lead to identical pictures. In the following text of this section, we will refer only to expressions for the eigenmodes obtained by the classical approach.

### v2

@@ II Model
III qualitative picture of magnon-polaritons
Before delving into detailed quantum and classical derivations, we present here a concise qualitative description of magnon-polariton physics in S/AF/S heterostructures. This overview highlights the key novel features of our proposal: the achievement of ultrastrong coupling at terahertz frequencies with exceptionally high cooperativity, the magnetic-field-tunable selectivity of the coupling, and the emergence of hybrid quasiparticles with remarkable properties, including a non-integer and $k$-dependent spin with high group velocity.
III.1 Formation mechanism
The core effect originates from the electromagnetic interaction between terahertz magnons in the antiferromagnetic (AF) layer and the photonic Swihart mode of the superconducting (S) resonator. The formation of magnon-polaritons requires three essential ingredients: the Swihart photon, the bare magnon, and their mutual dipolar interaction.
III.1.1 Swihart photon mode
The eigenfrequency of the Swihart mode in the superconducting resonator is given by [34]
$\Omega_{s}({\bm{k}})=\sqrt{\frac{d_{AF}}{\mu_{0}\varepsilon_{AF}(d_{AF}+\lambda_{\mathrm{eff}})}}|{\bm{k}}|,$ (1)
which vanishes in the limit $\lambda_{\mathrm{eff}}\to\infty$. Here $\mu_{0}$ is the vacuum magnetic permeability and $\varepsilon_{AF}$ is the dielectric constant of the antiferromagnet. Strictly speaking, a resonant mode persists even in a normal-metal environment, but the physics of magnon-polaritons is then strongly modified by dissipation [14]. For wavevectors aligned along the $z$-axis ($\bm{k}=k\bm{e}_{z}$), the magnetic field of the Swihart mode is polarized along the $y$-axis, as illustrated in Fig. 2(a). Further details on the Swihart mode and its quantization are provided in Sec. V.1.
Figure 2: Magnon-polaritons in S/AF/S heterostructures. (a) Magnetic field profile of the Swihart photon propagating in the superconducting resonator. (b) Magnetization precession of the antiferromagnetic magnon eigenmodes at zero and at a large applied magnetic field $\bm{H}_{0}=H_{0}\bm{e}_{z}$. (c) Dispersion of the hybridized eigenexcitations in the S/AF/S heterostructure: noninteracting case (photons and magnons decoupled, right panel) and interacting case (magnon-polariton formation, left panel). The averaged projection of the eigenexcitation spin $\langle S_{z}\rangle$ onto the $z$-axis is shown by color.
III.1.2 Magnons
The eigenmodes of bare magnons are derived in Sec. V.2. Neglecting demagnetization effects, the two eigenfrequencies correspond to the well-known antiferromagnetic resonance modes for an easy-axis antiferromagnet [35]:
where $\gamma$ is the electron gyromagnetic ratio, $J_{AF}$ is the antiferromagnetic exchange constant, $K$ is the easy-axis anisotropy constant, and $H_{0}$ is the external magnetic field along the Néel vector. The two modes are degenerate when $H_{0}=0$ due to the $PT$ symmetry.
In the S/AF/S geometry, the dynamic magnetization of the AF layer generates stray magnetic fields, which break the $PT$ symmetry and $U(1)$ symmetry around the Néel vector. It in turn induces Meissner screening currents in the adjacent superconducting layers. This current-field configuration strongly modifies the effective dipolar interaction within the AF medium. The demagnetization field $\bm{H}(\bm{r},t)=\tilde{\bm{H}}(x)e^{i\bm{k}\cdot\bm{\rho}-i\omega t}$ governs this interaction via the term
$\hat{H}_{\mathrm{dem}}=-\mu_{0}\sum_{i=1,2}\int d{\bm{r}}\hat{\bm{M}}_{i}({\bm{r}})\cdot{\bm{H}}(\bm{r},t).$ (3)
The demagnetization field, accounting for the influence of the superconducting environment, is calculated in Sec. IV and can be expressed in terms of a demagnetization tensor:
$\hat{H}=-\hat{N}\hat{M},\quad\hat{H}=\begin{pmatrix}\tilde{H}_{x}\\ \tilde{H}_{y}\end{pmatrix},\quad\hat{M}=\begin{pmatrix}\tilde{M}_{x}\\ \tilde{M}_{y}\end{pmatrix}$, (4)
with
$\hat{N}=\begin{pmatrix}1&0\\ 0&N_{y}(k)\end{pmatrix}$, (5)
where $N_{y}(k=0)\equiv N_{y0}=d_{AF}/(d_{AF}+\lambda_{\mathrm{eff}})$ and assumes a more complex form for $k\neq 0$. The bare magnons correspond to $k=0$, and their eigenfrequencies, including dipolar field effects, become
$\tilde{\omega}_{1,2}=\mu_{0}\gamma l+h_{+}\mp\sqrt{(\mu_{0}\gamma H_{0})^{2}+h_{-}^{2}},$ (6)
where $l=\sqrt{(K+J_{AF})^{2}-J^{2}_{AF}}$ and the terms $h_{\pm}$ arise from dipolar fields:
$h_{\pm}=\frac{1}{2}\mu_{0}\gamma M_{s}(1\pm N_{y0})\frac{K}{l}.$ (7)
Accordingly, even in the absence of the external magnetic field, the two antiferromagnetic modes are no longer degenerate due to the breaking of $PT$ symmetry.
III.1.3 Magnon-photon coupling
The coupling between magnons and photons occurs through the Zeeman interaction:
$\hat{H}_{\mathrm{int}}=-\mu_{0}\int dx\,d{\bm{\rho}}\,\hat{M}_{y}({\bm{r}})\hat{H}_{\mathrm{Sw}}({\bm{r}}).$ (8)
Since the magnetic field of the Swihart photon $\hat{H}_{\mathrm{Sw}}$ is along the $y$-axis, only the $y$-component of the magnetization couples to it. In quantum form, the interaction Hamiltonian takes the form (see Sec. V.3 for calculation details):
$\hat{H}_{\mathrm{int}}$ $=\hbar\sum_{\bm{k}}\Big(g_{1}({\bm{k}})\tilde{m}_{1,\bm{k}}(\hat{p}_{\bm{k}}^{\dagger}+\hat{p}_{-\bm{k}})$
$+g_{2}({\bm{k}})\tilde{m}_{2,\bm{k}}(\hat{p}_{\bm{k}}^{\dagger}+\hat{p}_{-\bm{k}})+\mathrm{H.c.}\Big),$ (9)
where $\hat{p}_{\bm{k}}$ are photon operators and $\tilde{m}_{1(2),\bm{k}}$ are magnon operators corresponding to the $\tilde{\omega}_{1,2}$ modes, respectively. The coupling constants $g_{1,2}$ are distinct. According to our calculations in Sec. V.3,
$g_{1}$ $\propto\frac{h_{-}}{\sqrt{h_{-}^{2}+\left(\mu_{0}\gamma H_{0}-\sqrt{(\mu_{0}\gamma H_{0})^{2}+h_{-}^{2}}\right)^{2}}}$
$\quad+\frac{h_{-}}{\sqrt{h_{-}^{2}+\left(\mu_{0}\gamma H_{0}+\sqrt{(\mu_{0}\gamma H_{0})^{2}+h_{-}^{2}}\right)^{2}}},$
$g_{2}$ $\propto\frac{\mu_{0}\gamma H_{0}-\sqrt{(\mu_{0}\gamma H_{0})^{2}+h_{-}^{2}}}{\sqrt{h_{-}^{2}+\left(\mu_{0}\gamma H_{0}-\sqrt{(\mu_{0}\gamma H_{0})^{2}+h_{-}^{2}}\right)^{2}}}$
$\quad+\frac{\mu_{0}\gamma H_{0}+\sqrt{(\mu_{0}\gamma H_{0})^{2}+h_{-}^{2}}}{\sqrt{h_{-}^{2}+\left(\mu_{0}\gamma H_{0}+\sqrt{(\mu_{0}\gamma H_{0})^{2}+h_{-}^{2}}\right)^{2}}}.$ (10)
III.2 Magnetic-field-controlled coupling selectivity
A distinctive feature of the S/AF/S system is the tunable selectivity of magnon-photon coupling by an external magnetic field $\mathbf{H}_{0}=H_{0}\mathbf{e}_{z}$, as suggested by Eqs. (10).
At zero field ($H_{0}=0$), the degeneracy of the two bulk AF magnon modes (see Eq. (2)) is lifted by the dipolar interaction (Eq. (6)). The resulting modes are nearly linearly polarized along the $x$ and $y$ axes, respectively, see Fig. 2(b). Only the mode with finite net magnetization along $y$ couples to the Swihart photon, forming a bright magnon-polariton with coupling constant $g_{1}\neq 0$. It is seen as upper and lower magnon-polariton branches, $\Omega_{2}$ and $\Omega_{3}$, respectively, in Fig. 2(c). The mode polarized along $x$ remains a dark mode, decoupled from the photon, consistent with $g_{2}=0$. It corresponds to the $\Omega_{1}$-mode in Fig. 2(c). At $H_{0}=0$, the bright mode ($\tilde{\omega}_{1}$) corresponds to sublattices precessing with opposite phases, producing a net $y$-magnetization, as shown in Fig. 2(b). The dark mode ($\tilde{\omega}_{2}$) involves precessions that cancel the net $y$-component. This intuitive classical picture is detailed in Sec. VI.2.
Under a finite magnetic field, the AF eigenmodes evolve from linear toward circular via elliptical polarization. Consequently, both magnon modes develop a finite $y$-component of magnetization and couple to the Swihart photon, see Fig. 2(b). The coupling strength for each mode varies smoothly with $H_{0}$, becoming equal only in the high-field limit ($\gamma\mu_{0}H_{0}\gg h_{-}$), as seen in Eqs. (10). This continuous tunability results from fully accounting for dipolar fields and Meissner screening, distinguishing our results from earlier works that neglected dipolar fields, where coupling was essentially binary: any nonzero field was effectively “strong” and gave equal coupling to both antiferromagnetic modes, with decoupling occurring only at $H_{0}=0$ [24]. The field scale $h_{-}$ governing this transition emerges from the dipolar interaction renormalized by the superconductors.
III.3 Ultrastrong coupling and high cooperativity
For typical parameters (e.g., using MnF2 as the prototypical AF and NbN as the superconductor), the coupling constant reaches $|g_{1}|\sim 100$ GHz. Given the terahertz-scale bare frequencies of AF magnons ($\tilde{\omega}\sim 1$ THz), the ratio $g/\tilde{\omega}\sim 0.1$ places the system at the boundary of the ultrastrong coupling regime. This coupling strength substantially exceeds values reported for AF magnon-polaritons in conventional terahertz cavities [24, 26, 27, 28, 29, 30, 31, 33] and in analogous superconductor/ferromagnet/superconductor heterostructures [17, 18, 19, 20]. Furthermore, the low dissipation in both the superconducting resonator and the high-quality AF insulator leads to an exceptionally high cooperativity $C=4g^{2}/(\kappa_{m}\kappa_{ph})\sim 10^{7}$, where $\kappa_{m}$ and $\kappa_{ph}$ are the magnon and photon decay rates, respectively, promising coherent quantum applications. See Sec. VI.1 for further details on the calculation of magnon and photon decay rates and cooperativity.
III.4 Properties of the hybrid magnon-polaritons.
The hybrid magnon-polariton branches (three in the general case; see Fig. 2(c)) inherit unique characteristics from their constituents.
Non-integer and $k$-dependent spin. The magnon-polaritons carry a non-integer average spin $\langle S_{z}\rangle<\hbar$, which depends on both the external field $H_{0}$ and the wavevector $k$. The $z$-component of spin is not conserved because terms in the quantum Hamiltonian originating from the dipolar interaction do not conserve the number of magnons with a definite spin projection onto the $z$-axis, as detailed in Sec. V.2. The average spin of bare magnon modes (before hybridization with the Swihart photon) is independent of $k$ but non-integer in general. However, it becomes $k$-dependent due to magnon-photon hybridization via electromagnetic interaction, as shown in Fig. 2(c). A striking prediction is the switching of the spin sign for the middle magnon-polariton branch upon varying $k$. A similar effect of spin non-conservation due to dipolar interactions was demonstrated in Ref. [36], but only for a ferrimagnet where bare magnon branches can intersect.
High group velocity. In the strong-mixing region where bare magnon and photon dispersions intersect, the group velocities of the magnon-polariton branches can reach a significant fraction of the speed of light ($v_{g}\lesssim c/4$). This vastly exceeds typical magnon group velocities and enables the prospect of ultrafast and superconductivity-tunable magnon transport in antiferromagnets.
In the following sections, we substantiate this qualitative picture with rigorous derivations. The calculation of modified dipolar fields in the superconducting environment is presented in Sec. IV. The quantum theory of magnon-photon coupling, leading to magnon-polariton dispersion and spin, is developed in Sec. V. For comparison and physical insight, Sec. VI re-derives the main results using the classical Landau-Lifshitz-Gilbert formalism, explicitly analyzing polarization states. Technical details are deferred to the appendices.
IV Dipolar fields
@@
where $\rho_{s}=\rho_{0}(1-(T/T_{c})^{4})$ and $\rho_{n}=\rho_{0}(T/T_{c})^{4}$ are superfluid and normal fluid densities, respectively. Equation (13) can then be rewritten as
@@
where $\sigma_{n}=\rho_{0}e^{2}\tau/m_{e}$ is the conductivity of normal metals, $\lambda_{L}=\sqrt{m_{e}/(\mu_{0}\rho_{s}e^{2})}$ is London’s penetration depth, $\rho_{0}$ is the electron density, $m_{e}$ is the electron mass, $e$ is the electron charge, $T_{c}$ is the critical temperature of the superconductor, $T$ is the temperature of the system, and $\tau$ is the relaxation time of electrons.
@@
On the other hand, after substituting explicit expressions for $\lambda_{L}$ and $\sigma_{n}$ Eq. (14) is transformed to
@@
$\mathrm{in~S1:}~~\tilde{\bm{E}}(x)=\bm{S}_{1}e^{iB_{S}(x-d_{AF})},$
$\mathrm{in~S2:}~~\tilde{\bm{E}}(x)=\bm{S}_{2}e^{-iB_{S}(x+d_{AF})},$
@@
where $A=\sqrt{k_{AF}^{2}-k^{2}}$, $B_{S}=\sqrt{k_{S}^{2}-k^{2}}$, and $\bm{S}_{1(2)}=(S_{1(2)x},S_{1(2)y},S_{1(2)z})^{T}$ and $\bm{A}^{\pm}=(A^{\pm}_{x},A^{\pm}_{y},A^{\pm}_{z})^{T}$ are vectors of unknown coefficients. Here we have used the assumption that the AF is thin enough, such that the magnetizations $\bm{M}_{1,2}$ can be considered as constant along the film normal $x$-axis throughout $2d_{AF}$. Also, for the considered parameters $B_{S}\approx i/\lambda_{\rm{eff}}$, which will be substituted in all the expressions containing $B_{S}$ in the following text.
@@
This dipolar field is responsible for the formation of magnon-polaritons with ultrastrong coupling between magnon and photon, which we address below by the quantum (Sec. V) and classical (Sec. VI) views, respectively.
V Magnon-polaritons: quantum view
@@ V.2 Uncoupled magnons
Then we introduce the operators of the spin density $\hat{\bm{S}}_{i}({\bm{r}})$ at the sublattice $i$, so that $\hat{\bm{M}}_{i}({\bm{r}})=-\gamma\hbar\hat{\bm{S}}_{i}({\bm{r}})$ and $M_{s}=\gamma\hbar S$, where $S$ is the total spin of the system. As the next step, we represent the spin-density operators $\hat{\bm{S}}_{i}({\bm{r}})$ in terms of the bosonic operators $\hat{a}_{i}({\bm{r}})$ via the Holstein-Primakoff transformation [42]:
@@
$+\frac{h_{-}}{\sqrt{h_{-}^{2}+\left(\mu_{0}\gamma H_{0}+\sqrt{(\mu_{0}\gamma H_{0})^{2}+h_{-}^{2}}\right)^{2}}},$
@@ V.4 Magnon-polaritons: dispersion
In the general case of arbitrary magnitudes of the applied magnetic fields, the eigenfrequencies of the magnon-polaritons are expressed by Eqs. (130) and are plotted in Fig. 3 for three typical values of the external magnetic field $H_{0}$: zero external field [Fig. 3(a)], moderate external field $\gamma\mu_{0}H_{0}\sim h_{-}$ [Fig. 3(b)], and strong external field $\gamma\mu_{0}H_{0}\gg h_{-}$ [Fig. 3(c)]. It is seen that at nonzero magnetic field both modes are coupled to the Swihart mode according to the fact that both coupling constants $g_{1}$ and $g_{2}$ are nonzero in this case. At large magnetic fields $\gamma\mu_{0}H_{0}\gg h_{-}$, both magnonic modes are equally coupled to the Swihart mode.
V.5 Magnon-polaritons: spin and group velocity
In addition to being at the boundary of the ultra-strong coupling, the complex magnon-polariton excitations have other important properties. First of all, they are bosonic quasiparticles with non-integer average spin $\langle S_{z}\rangle$ smaller than $\hbar$. The average spin of all magnon-polariton excitation branches is shown by the color of the lines in Fig. 3. The details of the calculation of the average spin are presented in the Appendix C.
[…]
@@ V.5 Magnon-polaritons: spin and group velocity
VI Magnon-polaritons: Classical view
VI.1 Classical derivation of the dispersion relations
Now we shall use an alternative classical approach to obtain the magnon-polariton eigenfrequencies and mode functions. The Landau-Lifshitz-Gilbert equation for the magnetizations of the two sublattices $\bm{M}_{1,2}$ takes the form:
$\dfrac{\partial\bm{M}_{i}}{\partial t}=-\gamma\mu_{0}\bm{M}_{i}\times\bm{H}_{i}{+\dfrac{\alpha}{M_{s}}\bm{M}_{i}\times\dfrac{\partial\bm{M}_{i}}{\partial t}},~~i\in\{1,2\},$ (70)
@@ VI.1 Classical derivation of the dispersion relations
is the total effective magnetic field acting on the magnetization of the sublattice $i$, $\overline{i}=2(1)$ if $i=1(2)$, and $\alpha$ is the Gilbert damping constant. Substituting Eq. (43) and a classical analogue of Eq. (46) into Eq. (71) and then linearizing Eq. (70) with respect to the magnon amplitudes, we obtain
$\left(\begin{array}[]{cccc}i\omega&0&-\gamma\mu_{0}(F_{+}+N_{y}M_{s}){+i\alpha\omega}&-\gamma\mu_{0}(J_{AF}+N_{y}M_{s})\\ 0&i\omega&\gamma\mu_{0}(J_{AF}+N_{y}M_{s})&\gamma\mu_{0}(F_{-}+N_{y}M_{s}){-i\alpha\omega}\\ \gamma\mu_{0}(F_{+}+N_{x}M_{s}){-i\alpha\omega}&\gamma\mu_{0}(J_{AF}+N_{x}M_{s})&i\omega&0\\ -\gamma\mu_{0}(J_{AF}+N_{x}M_{s})&-\gamma\mu_{0}(F_{-}+N_{x}M_{s}){+i\alpha\omega}&0&i\omega\end{array}\right)\left(\begin{array}[]{c}\tilde{M}_{1x}\\ \tilde{M}_{2x}\\ \tilde{M}_{1y}\\ \tilde{M}_{2y}\end{array}\right)=0,$ (80)
@@
are the eigenfrequencies in the case with no external magnetic field, $H_{0}=0$. In Eqs. (83a)-(83b), corrections on the order of $\alpha^{2}$ arising from Gilbert damping are neglected. The first-order corrections yield the decay rates of the modes $\omega_{01,2}$:
$\kappa_{01}$ $=\alpha\gamma\mu_{0}(J_{AF}+K+M_{s}),$ (84)
$\kappa_{02}$ $=\alpha\gamma\mu_{0}(J_{AF}+K+NM_{s}).$ (85)
Using the same parameters as in Fig. 3 and taking $\alpha\sim 10^{-4}$, the magnitude of these decay rates can be estimated as $\kappa_{01,2}\sim 1\ \mathrm{GHz}$.
This estimate enables us to evaluate the cooperativity of the magnon-polaritons, defined as $C=4g^{2}/(\kappa_{m}\kappa_{ph})$, where $\kappa_{m}$ and $\kappa_{ph}$ are the magnon and photon decay rates, respectively. For our S/AF/S heterostructure, with $\kappa_{ph}\approx 3\ \mathrm{MHz}$ [6, 19] and $\kappa_{m}\approx\kappa_{01,2}(k=0)\approx 1\ \mathrm{GHz}$, we obtain an exceptionally high cooperativity $C\approx 10^{7}$. This value significantly exceeds the cooperativities reported for AF-based light–matter systems [25, 27, 29, 32, 33], as well as those in S/F/S structures, where $C\sim 10^{4}$ has been reported [19].
Furthermore, the parameter $U=\sqrt{C\,(g/\omega)}$, which quantifies the degree of coherence in ultrastrongly coupled systems [48], also attains a very large value in our case, $U\sim 10^{3}$. This is comparable to the highest values reported for superconducting qubits.
As the demagnetization factor $N$ is expressed by Eq. (39) and depends on $\omega$ and $k$, Eq. (82) and Eq. (83b) are actually implicit equations for $\omega_{1,2}(k)$ and $\omega_{02}(k)$, respectively. Figure 4 shows the dispersion relations, calculated for the structure where the antiferromagnet is $\mathrm{MnF}_{2}$ with the same parameters addressed in Sec. V. Here we again consider the bulk volume configuration, i.e., $\theta=0$. We note that according to the quantum formalism, the coupling constants between magnons and Swihart photon vanish in the Damon-Eshbach configuration $\theta=\pi/2$. The dispersion relations calculated via the quantum approach from Eq. (130) and via the classical approach from Eq. (82) lead to identical pictures. In the following text of this section, we will refer only to expressions for the eigenmodes obtained by the classical approach.
VI.2 Polarization of the eigenmodes